OpenStudy (anonymous):

f(x) = 7x – 12, what is f^–1(x) ?

OpenStudy (anonymous):

y=7x-12 (rewrite without function notation) x=7y-12 (exchange the position of the x and y variables) x+12=7y (solve for y; add 12) y=(x+12)/7 (divide by 7) then give the inverse function a name \[f^{-1}(x)=\frac{1}{7}x+\frac{12}{7}\]

OpenStudy (anonymous):

so is the answer (x+12/7?

OpenStudy (anonymous):

(x+12)/7

OpenStudy (anonymous):

No, the answer is f^-1(x) = (1/7)x + 12/7

OpenStudy (anonymous):

one interesting characteristic re. a function and its inverse is that their graphs are symmetric about the line y=x (see attached file)